A man swimming downstream overcome a float at a point M.After travelling distance D he turned back and passed the float at a distance of D/2 from the point M, then the ratio of speed of swimmer with respect to still water to the speed of the river will be???

Grade:11

1 Answers

SAHIL
7 years ago
SWIMMING DOWNSTREAM WITH DISTANCE D WITH SPEED= S+R
SWIMMING DOWNSTREAM WITH DISTANCE D/2 WITH SPEED= S-R
TOTAL TIME=D/S+R + D/2(S-R)=Dr/2
solving this we get
s/r=3
please post if you donot understand anywhere
regards sahil